OpenStudy (tay1lynn):

simplify 2x +64/ 3over 4

OpenStudy (itz_sid):

\[\frac{ 2x+64 }{ \frac{ 3 }{ 4 } }\] Is that your question?

OpenStudy (tay1lynn):

yes

OpenStudy (itz_sid):

\[\frac{ 2x +64 }{ 1 } \div \frac{ 3 }{ 4 }\] Okay, well you first start off by taking the reciprocal of the denominator. When you take the reciprocal, the division sign turns into multiplication. \[\frac{ 2x+64 }{ 1 }\times \frac{ 4 }{ 3 }\] So you get...\[\frac{ 4(2x+64) }{ 3 }\]

OpenStudy (itz_sid):

Then you factor the 4 over, can you do that yourself?

OpenStudy (tay1lynn):

yes thank you